OpenStudy (anonymous):

What is the area of a circle with diameter 8 inches? choices: 4 pi in^2 8 pi in^2 16 pi in^2 64 pi in^2

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow Area(Circle) = \pi({d \over 2})^2 }\)

OpenStudy (callisto):

Now, d=8

Parth (parthkohli):

Or, you could say: \(\Large \color{MidnightBlue}{\Rightarrow Area(Circle) = \pi r^2 }\)

OpenStudy (anonymous):

is it 16 pi in^2?

Parth (parthkohli):

Since d/2 = r, we can use both

OpenStudy (callisto):

@HiImOlivia Yes

Parth (parthkohli):

Yes..